# GMP-ATL: Gender-augmented Multi-scale Pseudo-label Enhanced Adaptive Transfer Learning for Speech Emotion Recognition via HuBERT

### v1

@@ 5 CONCLUSIONS
In this work, we propose an effective transfer learning-based SER framework GMP-ATL, optimizing the pre-trained HuBERT model to discern emotions by incorporating more reasonable emotional pseudo-labels. Extensive experiments on the IEMOCAP corpus indicate that our proposed GMP-ATL framework not only outperforms SOTA unimodal SER methods but achieves competitive performance compared to multimodal SER approaches. We demonstrate that incorporating reasonable frame-level gender-augmented multi-scale pseudo-labels can effectively enhance the recognition performance of the proposed workflow. Furthermore, we observe that within our proposed framework, the features extracted from the third-to-last layer of the pre-trained HuBERT model seem to encompass more emotion-related information, a phenomena that may be applicable to other speech attributes as well.

### v2

@@ 5 CONCLUSIONS
In this work, we propose GMP-TL, an effective GMPs-based transfer learning framework for SER. Initially, GMP-TL employs a pre-trained HuBERT model, implemented with multi-task learning and multi-scale k-means clustering strategies, to obtain high-quality frame-level GMPs. Moreover, an efficient two-stage hybrid loss based fine-tuning approach is introduced to further optimize the GMP-TL by comprehensively leveraging frame-level GMPs and utterance-level emotion labels. Experiments on IEMOCAP show that our GMP-TL not only attains superior performance compared to SOTA unimodal SER methods but achieves competitive results compared to multimodal SER approaches, demonstrating the effectiveness of the proposed approach. In future work, we aim to develop more accurate frame-level emotional pseudo-labels by integrating additional relevant speech attributes, thereby further enhancing the proposed GMP-TL framework.